OpenStudy (anonymous):

HELP!!!!! If the ratio of the sides of two similar prisms is 3:5 and the volume of the first prism is 54, what is the volume of the second prism? A. 11.7 B. 32.4 C. 90 D. 250

OpenStudy (anonymous):

@johnweldon1993 @mathmale

OpenStudy (anonymous):

If the ratio of the sides is 3:5...the ratio of the volumes will be 3^3:5^3 or 27:125 Does that help? Can you take it from there?

OpenStudy (anonymous):

can work that out

OpenStudy (anonymous):

can u work that out

OpenStudy (anonymous):

not sure how to solve

OpenStudy (johnweldon1993):

Well it's true...the ratio would be the cubes of the side ratio...so 3^3 : 5^3 = 27 : 125 This would be the volume ratio... If the first prism is the smaller one...we can denote that as the "27" side... so the volume of the smaller is 54 \[\large \frac{54}{x} = \frac{27}{125}\] Now you can solve for 'x'

OpenStudy (anonymous):

I got 62.5

OpenStudy (johnweldon1993):

Try it again...cross multiply first...what do you get?

OpenStudy (anonymous):

oh cross mulitiply is 250 yay!! thanks